OpenStudy (anonymous):

2. Solve the equation. Round the answer to the nearest tenth. 1.04x + 3.14 = 7.98 (Points : 1)

OpenStudy (anonymous):

4.7

OpenStudy (pdoctor):

Subtract 3.14 from both sides, then divide both sides by 1.04 \[(1.04x + 3.14)-3.14=7.98-3.14\] \[1.04x=4.84\] \[(1.04x)/1.04 = 4.84/1.04\] \[x = 4.65384615\] \[x=4.7\]

OpenStudy (anonymous):

Thank Yall Very Much(: